OpenStudy (anonymous):

Identify the factors of x^2 - 8x - 20. (x + 4)(x - 5) (x + 5)(x - 4) (x - 10)(x + 2) (x - 2)(x + 10)

OpenStudy (anonymous):

Factor the following: x^2-8 x-20 Factor x^2-8 x-20 by finding two numbers whose product is -20 and whose sum is -8. The factors of -20 that sum to -8 are 2 and -10. So, x^2-8 x-20 = (x+2) (x-10): so the answer is | (x+2) (x-10)

OpenStudy (haseeb96):

this is the middle term break method x^2+2x -10x -20 =0 x(x+2) -10(x+2)=0 (x-10)(x+2) is the correct answer

OpenStudy (anonymous):

Thank you!

OpenStudy (haseeb96):

most welcome my dear